OpenStudy (anonymous):

An archway will be constructed over a walkway. A piece of wood will need to be curved to match a parabola. Explain to Maurice how to find the equation of the parabola given the focal point and the directrix.

OpenStudy (anonymous):

draw a perpendicular line from the directrix passing through the focus, this will be the line of symmetry. The vertex(h, k) will be located on the line half way between the focus and directrix. The distance from the focus to the vertex is called the focal length, call it a. The then equation is (x - h)^2 = 4a(y - k) the equation can be manipulated to y = 1/4a(x - h)^2 + k hope it helps

OpenStudy (anonymous):

ahahha

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

lets say the required equation of parabola is \(\large y = a(x-h)^2 + k\) where vertex = \((h, k)\)

OpenStudy (anonymous):

Ok. That makes sense so far

ganeshie8 (ganeshie8):

knw how to get the `vertex point` form `Focus` and `directrix` ?

OpenStudy (anonymous):

honestly no, the lesson didn't talk much about directx and don't even remember seeing focus

ganeshie8 (ganeshie8):